OpenStudy (inspectorjoe):

(2x+3y+4)(2x+3y-4) My answer is : 4x^2+12xy+9y2-20y. ?

OpenStudy (lynfran):

when we use foil we get 2x*2x+2x*3y-(2x*4)+3y*2x+3y*3y-(3y*4)+4*2x+4*3y-(4*4) 4x^2+6xy-8x+6xy+9y^2-12y+8x+12y-16 combine like term add/subtract...

OpenStudy (lynfran):

final answer 4x^2+12xy+9y^2-16

OpenStudy (ddcamp):

There's a nifty trick for expanding certain polynomials. (a+b)(a-b) = a² - b² Here, let a = 2x + 3y, and b = 4: (2x + 3y)² - 4² is a fair bit simpler to expand, isn't it?

OpenStudy (lynfran):

yep indeed @DDCamp its much easier

OpenStudy (inspectorjoe):

But what happen with the 8x? I'm missing something

OpenStudy (inspectorjoe):

Oh I fo be it and just at got

OpenStudy (lynfran):

the 8x cancels out

OpenStudy (inspectorjoe):

Yes

OpenStudy (lynfran):

im telling u that not asking you...lol

OpenStudy (inspectorjoe):

Lynn here is my next problem and what I came up with.. Let me know if can help.

OpenStudy (inspectorjoe):

Sorry I was answering lol

OpenStudy (lynfran):

ok

OpenStudy (inspectorjoe):

(5x+2y+3)(5x+2y-3) This is what I came up with: 25x^2+20xy-15x+4y^2-6y?

OpenStudy (inspectorjoe):

Wait wait I gave you my incorrect answer I will give you my results right now

OpenStudy (inspectorjoe):

My correct answer is: 25x^2+20xy+4y^2-9

OpenStudy (inspectorjoe):

The -15 clear with the positive 15..

OpenStudy (inspectorjoe):

InspectorJoe 0 My correct answer is: 25x^2+20xy+4y^2-9

OpenStudy (lynfran):

yep ur correct

OpenStudy (inspectorjoe):

YES

OpenStudy (inspectorjoe):

YEAH

OpenStudy (inspectorjoe):

Lol

OpenStudy (inspectorjoe):

Oh yes